Consider the function y=tan(x)

(a) Show that the first two non-zero terms in the Maclaurin series of y are x+1/3x3...

(b) Use the first two terms of the Maclaurin series of y to estimate tan (1/3).



1
Expert's answer
2021-09-22T17:51:54-0400


Maclaurin series of y(x) is

"y(x)=y(0)+y'(0)\cdot x+y''(0)\cdot x^2/2+y'''(0)\cdot x^3/6+o(x^3)"

We have

y(0)=tan(0)=0;

y'(x)=tan'(x)="\frac{1}{cos^2(x)}" ; y'(0)="\frac{1}{cos^2(0)}=\frac{1}{1}=1;"

y''(x)="\left( \frac{1}{cos^2(x)}\right)'=\frac{1'\cdot cos^2(x)-1\cdot cos^2(x))'}{cos^4(x)}=\\"

="\frac{2\cdot cos(x)\cdot sin(x) }{cos^4(x)}=\frac{2\cdot sin(x)}{cos^3(x)};"

"y''(0)=\frac{1\cdot sin(0)}{cos^3(0)}=\frac{2\cdot 0}{1}=0;"

y'''(x)="\left (\frac{2\cdot sin(x)}{cos^3(x)} \right)'=\frac
{2\cdot (sin(x))'\cdot cos^3(x)-2\cdot sin(x)\cdot(cos^3(x))'}
{cos^6(x)}=\\"

"=\frac{2\cdot cos^4(x)+6\cdot sin^2(x)\cdot cos^2(x)}{cos^6(x)}=
\frac{2}{cos^2(x)}-\frac{6\cdot tan^2(x)}{cos^2(x)};"

y'''(0)="\frac{2}{cos^2(0)}-\frac{6\cdot tan^2(0)}{cos^2(0)}=\frac{2}{1}-\frac{6\cdot 0}{1}=2"

Thus we have

y(x)=tan(x)="0+1\cdot x+0\cdot x^2/2+2\cdot x^3/6+o(x^3)=x+\frac{x^3}{3}+o(x^3)"

(b) Use the first two terms of the Maclaurin series of y to estimate tan (1/3)

tan(1/3)="\frac{1}{3}+\frac{\left( \frac{1}{3} \right)^3}{3}=\frac{1}{3}+\frac{1}{81}=
\frac{28}{81}=0.345679" -approx value

Exat value with 6 digits:


Thus error rate=0.0006 "\approx 0,001" by order